A 3.1 kg rifle fires an 11 g bullet with a velocity of 575m/s. What is the recoil velocity of the rifle? If the shooter holds the rifle firmly against his shoulder what will be the recoil velocity? Assume the mass of the shooter is 45.0kg


1
Expert's answer
2021-10-13T15:42:37-0400
"0.011(575)=3.1v_r\\v_r=2.04\frac{m}{s}"

"0.011(575)=(3.1+45)v_r'\\v_r'=0.131\frac{m}{s}"